OpenStudy (anonymous):

Find the inverse of the function f(x) = 4x − 28

OpenStudy (anonymous):

wouldnt the inverse just be the same function with an inverse of slope?

OpenStudy (anonymous):

let y=f(x) so we can write it as: y = 4x - 28 now switch the x and the y... then solve for y...

OpenStudy (anonymous):

that will give you the inverse of f...

OpenStudy (anonymous):

I guess im right too. My way is easier just inverse the slope The slope is 4 the inverse is -1/4

OpenStudy (anonymous):

replace and you have the answer.

OpenStudy (anonymous):

well not not really

OpenStudy (anonymous):

@romero what about the intercept?

OpenStudy (anonymous):

it is not right

OpenStudy (anonymous):

Prove it

OpenStudy (anonymous):

\[f(x) = 4x − 28\] function says a) multiply by 4 b) subtract 28 inverse says a) add 28 b) divide by 4 write is as \[f^{-1}(x)=\frac{x+28}{4}=\frac{1}{4}x+7\] a line with slope \(\frac{1}{4}\) not \(-\frac{1}{4}\)

OpenStudy (anonymous):

Gotcha bro I was looking for the perpendicular line.

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

y = 4x - 28 y - 28 = 4x (y - 28) / 4 = x now substitute place of x and y y^-1 = (x-28) /4 = x/4 - 7